# MFIT : Multi-FIdelity Thermal Modeling for 2.5D and 3D Multi-Chiplet Architectures

### v2

@@ IV-C FEM to Thermal RC models
This section describes the process of constructing a thermal RC model from the geometry of a given package. MFIT applies this technique to both 2.D and 3D systems demonstrating the flexibility of the proposed methodology to different packaging technologies. However, this process can easily be applied to any package.
The package is first divided into horizontal layers, with the slicing process starting at the bottom substrate layer and ending at the top lid layer. Depending on the package design, each layer may be composed of uniform material or various material blocks. This flexibility enables the thermal RC model to simulate packages with heterogeneous designs where different chiplets are manufactured with various technologies, resulting in different material parameters in the same layer. For layers with uniform material properties, it is divided into a 2D grid of nodes. When a layer contains different material blocks, each block can be divided into grids with different levels of granularity. a 3D network of thermal nodes that discretizes the chiplet geometry in space.
@@
$G_{x}=\frac{k_{x}\cdot l_{y}\cdot l_{z}}{l_{x}},~{}G_{y}=\frac{k_{y}\cdot l_{x}\cdot l_{z}}{l_{y}},~{}G_{z}=\frac{k_{z}\cdot l_{x}\cdot l_{y}}{l_{z}}$ (4)
@@
Heat is dissipated from the package primarily through a heatsink which is simulated using a convective heat transfer coefficient as detailed in Section IV-B3. MFIT assumes forced convection is applied to the heatsink while passive convection occurs on the other external boundaries of the package. Consequently, convective conductance ($G_{conv}=h_{eq}\cdot l_{x}\cdot l_{y}$) is incorporated into the nodes of the top and bottom layers.
The conductance between neighboring nodes $i$ and $j$ ($G_{ij}$) of the same layers is determined by lateral conductances ($G_{x}$ and $G_{y}$). Unlike existing thermal RC models [44, 16, 31], our thermal RC model allows non-uniform grid sizes for different layers and blocks, so a node in one layer can be connected to multiple nodes from adjacent layers. Thus, vertical conductances between nodes of different layers are calculated from $G_{z}$, considering the overlap in the x-y plane. Once this RC network is established, we can formulate an ODE based on Kirchhoff’s current law for a node $i$ as:
$C_{i}\frac{dT_{i}}{dt}=\textstyle\sum_{j=1}^{N}(G_{ij})(T_{j}-T_{i})+\dot{q}_{i}$ (5)
@@
$\mathbf{C}\times\dot{\mathbf{T}}=\mathbf{G}\times\mathbf{T}+\dot{\mathbf{q}}$ (6)
@@
$\mathbf{G}=\begin{bmatrix}-\sum_{j=1}^{N}G_{1j}&G_{12}&\cdots&G_{1N}\\ G_{21}&-\sum_{j=1}^{N}G_{2j}&\cdots&G_{2N}\\ \vdots&\vdots&\ddots&\vdots\\ G_{N1}&G_{N2}&\cdots&-\sum_{j=1}^{N}G_{Nj}\end{bmatrix}$ (7)
@@
MFIT employs the highly adaptive solver LSODA [45] to solve this system of ODEs. LSODA is designed to handle both stiff and non-stiff systems efficiently. It dynamically switches between different numerical integration methods depending on the characteristics of the ODE system being solved. This switching capability is particularly useful for thermal ODEs, as the equations’ stiffness can vary over time depending on the power consumption. It is worth noting that the matrices representing the system are highly sparse since each node is connected to only a few neighboring nodes. MFIT leverages this sparsity to accelerate the solver’s execution time. Finally, we fine-tune the capacitance values of each layer utilizing FEM results as a reference to improve the accuracy of our model.
IV-D Thermal RC to Discrete State Space models
The thermal RC model can be discretized in the time domain to further reduce the execution time of the model with no cost of accuracy. The resulting DSS model’s limitation is its dependence on an underlying continuous time model. It cannot be constructed directly without a thermal RC model as an intermediate step or system identification and measurement data. Additionally, a DSS model is specific to the geometry, materials, and sampling period used during the creation of the thermal RC model and the later discretization process. Therefore, the DSS model must be reconstructed if any design parameter changes. Only an existing RC model and a set time step are required to create a DSS model, with no direct information from the previous FEM model being needed.
Integrating the continuous time system, as given in Equation 6 for $t=t+T_{s}$ results in [46]:
@@ IV-D Thermal RC to Discrete State Space models
Assuming constant heat generation during the sampling period $T_{s}$, the system can be discretized from continuous time variable $t$ to discrete steps $k$ as:
$\mathbf{T}[k+1]=\mathbf{A}\times\mathbf{T}[k]+\mathbf{B}\times\dot{\mathbf{q}}[k]$ (9)
where $\mathbf{A}$ and $\mathbf{B}$ are the state and input matrices. Equation 9 represents the discrete-time equivalent of the continuous-time thermal RC model (shown in Equation 6). MFIT uses the zero-order hold (ZOH) method for the discretization process. When power is provided as discrete inputs at each sampling period, ZOH provides an exact match to the continuous time model. $T_{s}$ can be determined for discretization as a function of input power consumption and system dynamics.
DSS model consists only of multiply-accumulate operations, allowing for extremely fast operation, as shown in Section V. The discretization process is also nearly instantaneous, allowing for rapid DSS model creation when a thermal RC model is available.
V Experimental Results
V-A Experimental Setup
@@
TABLE V: Specifications of simulated systems in this work
@@
Package
Thickness (mm)
1.855 1.855 1.855 2.255
Package Length
and Width (mm)
15.5 21.5 27.5 15.5
Package Top
Area (mm2)
240.25 462.25 756.25 240.25
Package Volume (mm3) 445.66 857.47 1402.84 534.55
@@
Individual Chiplet
Power (W)
3 3 3 1.2
Total System
Power (W)
48 108 192 57.6
Total System Power
per lid area (W/mm2)
0.199 0.233 0.253 0.239
@@
Maximum Chiplet
Temperature (C)
118.25 129.75 164.03 128.65
V-A1 Package Overview
@@
Both the 2.5D and 3D systems utilize a silicon interposer with chiplets placed upon it. The interposer is connected to the underlying substrate using C4 bumps. Copper wires embedded in the interposer are used to connect neighboring chiplets. In both our 2.5D and 3D systems, each chiplet area is considered to be $2.25mm^{2}$, consistent with prior studies [19, 20, 5]. Each chiplet consists of multiple blocks. Each of these blocks corresponds to a component such as a computational tile or a router used for inter-chiplet communication, as detailed in Figure 1(a). Each of these blocks which make up the chiplet has an individual power profile. This means that different power profiles can be applied to every computational tile and router port in each chiplet. In our experimentation, different levels of detail are applied to the chiplets in the 2.5D and 3D system, as described in the following sections.
@@
2.5D System Specifics: The target 2.5D system consists of a grid of chiplets integrated on an interposer, as illustrated in Figure 1(a). Each chiplet is connected directly to the interposer via $\mu$-bumps surrounded by a capillary underfill material. The physical dimensions of the router ports are compatible with Universal Chiplet Interconnect Express (UCIe) specification [47]. The entire package is covered by a copper lid, which contacts each chiplet through a thermal interface material (TIM).
@@
3D System Specifics: In the 3D system, three stacked chiplets are placed in a 4x4 grid with equal spacing, consistent with [9]. The bottom chiplets are connected to the interposer through the $\mu$-bumps surrounded by a capillary underfill material, as detailed in Figure 1(b). The lid contacts only the top chiplet layer through a thermal interface material.
V-B Thermal RC Model Configuration
@@
The number of nodes in the thermal RC network determines the model complexity, runtime, and granularity at which temperature can be observed in the model. A higher node density is used in chiplets to optimize this trade-off, while fewer nodes are used in non-chiplet components, such as the interposer, lid, substrate, and so on. Each chiplet is divided into four equal quadrants. One node is placed within each quadrant to allow for granular temperature monitoring across each chiplet. For non-chiplet layers in each model, an alternate node density is used for each model as described below. This easily configurable non-uniform node density enables higher thermal resolution in critical parts such as chiplets while decreasing the runtime with lower resolution in less critical structures such as the substrate and lid. The DSS models in our experimentation are created by discretizing the thermal RC models with $T_{s}=0.01s$ sampling period. The sampling time can be chosen as a function of the application requirements.
@@
Figure 8: Execution times (in log scale) of the proposed thermal models and HotSpot [15] for 16, 36, 64 - 2.5D integrated and 16x3 - 3D integrated chiplet systems.
3D Thermal RC Model Specifics: The node densities in the 3D system are adjusted similar to the 2.5D models. The layers that contain chiplets use a 8$\times$8 grid, implying 4 nodes per chiplet. All other layers have a 4$\times$4 grid, leading to a lower node density. The entire 3D system consists of 48 chiplets in total. There are a total of 192 nodes counting all nodes within chiplets.
V-B1 Input Workloads and Power Consumption
Identical workloads are considered for the 2.5D and 3D systems, with differences in chiplet power density detailed in the following subsections. The target chiplet systems are analyzed under one synthetic (WL1) and five real AI/ML application workloads (WL2-WL6). The synthetic workload starts with a stress test that applies the maximum power to all chiplets to increase temperature beyond 100C. Then, a pseudo-random bit sequence (PRBS) is applied to each chiplet to emulate a wide range of dynamic variations. Finally, all chiplets are turned off to let the temperature return to the ambient state, as depicted in Figure 10. Besides testing transient and steady-state behaviors, this power profile helps us to tune the thermal RC model.
The remaining scenarios consider processing-in-memory (PIM)-based chiplets for accelerating ML workloads. The computational platform is resistive random access memory (ReRAM) based chiplets commonly used in literature [20, 4, 48]. We select this configuration due to it’s ability to efficiently implement matrix-vector multiplication, which is the predominant operation in any CNN workload. Each workload consists of a series of deep neural networks (DNNs) which run in series on the system. The workloads are listed in Table VI. The neural networks (NN) in these workloads consist of several networks such as ResNets, DenseNets, and VGG networks. For example, WL1 contains 16 ResNet34’s, followed by one VGG19, then 5 ResNet50’s, and so on. Each workload contains from 20 to 40 individual networks. Workloads are mapped to the system as computing resources become available, meaning a new NN is mapped to chiplets when it completes the execution of a previous NN. Consequently, these workloads consist of NNs ranging from small NNs like ResNet18, which can be mapped to a single chiplet, to larger NNs such as DenseNet169, which are spread across multiple chiplets.
TABLE VI: Descriptions of the workloads used in this work. (C) denotes CIFAR100 dataset, (I) denotes ImageNet dataset.
@@ V-B1 Input Workloads and Power Consumption
After mapping the group of NNs to the target chiplet-based systems, the chiplet power consumption is estimated in two parts, communication and computation. We estimate computation power through NeuroSim and interconnection network power using BookSim [49, 50]. We use running average power throughout the workload execution (40-55 seconds), consistent with power measuring tools such as Intel RAPL [51] and pyNVML [52].
@@
Differences in 2.5D and 3D system chiplet power: The experiments use different hardware parameters, such as voltage and frequency, for the 2.5D and 3D systems. This results in lower per-chiplet power consumption in the 3D system of 1.2W as compared to 3W for the 2.5D system, detailed under the Power section of Table  V. Using these parameters, the total system power per lid area of the 3D system is between that of the 36 and 64 chiplet 2.5D system. This level means that the temperature of individual chiplets should be roughly equivalent between these systems, which is confirmed in Figure 10.
V-B2 HotSpot [15] configuration:

### v4

@@ 4.3. FEM to Thermal RC models
The package is first divided into horizontal layers, with the slicing process starting at the bottom substrate layer and ending at the top lid layer. Depending on the package design, each layer may be composed of a uniform material or various material blocks, resulting in either homogeneous or non-homogeneous layers. This flexibility enables the thermal RC model to simulate packages with heterogeneous designs where different chiplets are manufactured with various technologies, resulting in different material parameters in the same layer. Layers with uniform material properties are divided into a 2D grid of nodes, where the grid granularity can vary between layers. For non-homogeneous layers with different material blocks, each block can have a distinct grid granularity, resulting in a non-uniform grid that connects the entire package as a 3D network of thermal nodes, discretizing the chiplet geometry in space.
@@
(4) $G_{x}=\frac{k_{x}\cdot l_{y}\cdot l_{z}}{l_{x}},~{}G_{y}=\frac{k_{y}\cdot l_{x}\cdot l_{z}}{l_{y}},~{}G_{z}=\frac{k_{z}\cdot l_{x}\cdot l_{y}}{l_{z}}$
@@
Heat is dissipated from the package primarily through a heatsink which is simulated using a convective heat transfer coefficient as detailed in Section 4.2.4. MFIT assumes forced convection is applied to the heatsink while passive convection occurs on the other external boundaries of the package. Consequently, convective conductance ($G_{conv}=h_{eq}\cdot l_{x}\cdot l_{y}$) is incorporated into the nodes of the top and bottom layers.
The conductance between neighboring nodes $i$ and $j$ ($G_{ij}$) of the same layers is determined by lateral conductance ($G_{x}$ and $G_{y}$). Since our thermal RC model allows non-uniform grid sizes for different layers and blocks, a node in one layer can be connected to multiple nodes from adjacent layers. Thus, vertical conductance between nodes of different layers is calculated from $G_{z}$, considering the overlap in the x-y plane. Once this RC network is established, we can formulate an ODE based on Kirchhoff’s current law for a node $i$ as:
(5) $C_{i}\frac{dT_{i}}{dt}=\textstyle\sum_{j=1}^{N}(G_{ij})(T_{j}-T_{i})+\dot{q}_{i}$
@@
(6) $\mathbf{C}\times\dot{\mathbf{T}}=\mathbf{G}\times\mathbf{T}+\dot{\mathbf{q}}$
@@
(7) $\mathbf{G}=\begin{bmatrix}-\sum_{j=1}^{N}G_{1j}&G_{12}&\cdots&G_{1N}\\ G_{21}&-\sum_{j=1}^{N}G_{2j}&\cdots&G_{2N}\\ \vdots&\vdots&\ddots&\vdots\\ G_{N1}&G_{N2}&\cdots&-\sum_{j=1}^{N}G_{Nj}\end{bmatrix}$
@@
MFIT employs the backward Euler method to solve this system of ODEs, as its implicit formulation ensures stability when handling the stiff equations typical of thermal modeling. Given that the system’s matrices are highly sparse (each node connects to only a few neighbors), MFIT leverages this sparsity by integrating SuperLU (Li, 2005) as a sparse linear solver. Additionally, we incorporate the BLAS library (Ref, [n. d.]) alongside SuperLU to further accelerate computations. To balance flexibility and performance, we implement the front-end in Python to define the geometry and ODE Equations, while the core computations rely on optimized C-based implementations of SuperLU.
Capacitance Tuning: While the thermal RC method provides very accurate steady-state temperature predictions, transient temperature calculations can exhibit relatively higher errors. This is a limitation of existing approaches as discussed in the experimental evaluations. To mitigate this issue, we fine-tune the capacitance values of each layer in the system using FEM results as a reference. Specifically, we introduce a scalar multiplier for each layer’s capacitance and optimize these values using the nonlinear optimizer Nelder-Mead (Barton and Ivey Jr, 1996). This tuning process is performed on a small-scale system and is based on the number of layers and their material properties rather than chiplet placement or geometry. Once optimized, the same capacitance parameters are applied to larger system sizes for more accurate thermal modeling without increasing grid complexity, as we demonstrate in Section 5.4. Since most advanced packaging technologies rely on similar materials, re-tuning is rarely required. Tuning is only necessary when the number of layers or material properties changes, in which case fine-tuning capacitance values on a small-scale system ensures accurate temperature calculation. In this work, we tune two representative small-scale systems, one for 2.5D systems and one for 3D system. The capacitance values of these representative systems are applied to the larger systems shown in the experimental results.
4.4. Thermal RC to Discrete State Space models
The thermal RC model can be discretized in the time domain to further reduce the execution time of the model with no cost of accuracy. The discretization of a thermal RC model is completed by automated tools and takes only milliseconds to run. The DSS model can be used for dynamic thermal management applications where the geometry is fixed, and faster thermal prediction is more important. However, a limitation of the DSS model is its dependence on an underlying continuous time model. It cannot be constructed directly without a thermal RC model as an intermediate step or system identification and measurement data. Additionally, a DSS model is specific to the geometry, materials, and sampling period used during the creation of the thermal RC model and the later discretization process. Therefore, the DSS model must be reconstructed if any design parameter changes. Only an existing RC model and a set time step are required to create a DSS model, with no direct information from the previous FEM model being needed.
Rearranging the Equation 6 into a state-space representation:
(8) $\dot{\mathbf{T}}=\mathbf{A}\mathbf{T}+\mathbf{B}\dot{\mathbf{q}},\quad\text{where }\mathbf{A}=\mathbf{C}^{-1}\mathbf{G},\quad\mathbf{B}=\mathbf{C}^{-1}.$
The time evolution of this system over an interval $t\in[kT_{s},(k+1)T_{s}]$, where $T_{s}$ denotes the sampling period, is determined by solving the differential equation.
In the absence of power ($\dot{\mathbf{q}}=0$), the homogeneous solution is given as:
(9) $\mathbf{T}(t)=e^{\mathbf{A}(t-kT_{s})}\mathbf{T}(kT_{s})$
Assuming $\dot{\mathbf{q}}$ is present and remains constant over the interval under a zero-order hold (ZOH) approximation, the particular solution is given as:
@@ 4.4. Thermal RC to Discrete State Space models
Combining Equations 9 and  10:
[…]
Evaluating at $t=(k+1)T_{s}$, the system is discretized as:
[…]
Then, the discrete-time system matrices is defined as:
(13) $\mathbf{A_{d}}=e^{\mathbf{A}T_{s}},\quad\mathbf{B_{d}}=\mathbf{A}^{-1}(\mathbf{A_{d}}-\mathbf{I})\mathbf{B}$
the resulting discrete-time state-space equation is given by:
(14) $\mathbf{T}[k+1]=\mathbf{A_{d}}\mathbf{T}[k]+\mathbf{B_{d}}\dot{\mathbf{q}}[k]$
Equation 14 represents the discrete-time equivalent of the continuous-time thermal RC model (shown in Equation 6). MFIT uses the zero-order hold (ZOH) method as given in Equation 10 for the discretization process. When power is provided as discrete inputs at each sampling period, ZOH provides an exact match to the continuous time model. $T_{s}$ can be determined for discretization as a function of input power consumption and system dynamics.
The DSS model consists only of multiply-accumulate operations, allowing for extremely fast operation, as shown in Section 5. The discretization process has only a one-time cost and is also nearly instantaneous, allowing for rapid DSS model creation when a thermal RC model is available.
5. Experimental Results
5.1. Experimental Setup
@@
Table 7. Specifications of simulated systems in this work.
@@
Package Thickness (mm) 1.855 1.855 1.855 2.105
Package Length and Width (mm) 15.5 21.5 27.5 15.5
Package Top Area (mm2) 240.25 462.25 756.25 240.25
Package Volume (mm3) 445.66 857.47 1402.84 505.72
@@
Individual Chiplet Maximum Power (W) 3 3 3 1.2
Total System Maximum Power (W) 48 108 192 57.6
Total System Maximum Power per lid area (W/mm2) 0.199 0.233 0.253 0.239
@@
Maximum Chiplet Temperature (C) 118.25 129.75 164.03 142.01
5.1.1. Package Overview
@@
Both the 2.5D and 3D systems utilize a silicon interposer with chiplets placed upon it. The interposer is connected to the underlying substrate using C4 bumps. A copper lid covers the chiplets and acts as a heat spreader. The lid and substrate form the outer bounds of the package. Dimensions of the package for each system configuration are found in Table 7. Within the package, copper wires embedded in the interposer are used to connect neighboring chiplets. In both our 2.5D and 3D systems, each chiplet area is considered to be $2.25mm^{2}$, consistent with prior studies (Krishnan et al., 2021; Sharma et al., 2022a, 2023). Each chiplet consists of multiple blocks. Each of these blocks corresponds to a component, such as a computational tile or a router used for inter-chiplet communication, as detailed in Fig. 1(a). Each of these blocks which make up the chiplet has an individual power profile. This means that different power profiles can be applied to every computational tile and router port in each chiplet. In our experimentation, different levels of detail are applied to the chiplets in the 2.5D and 3D systems as described in the following sections.
@@
2.5D System Specifics: The target 2.5D system consists of a grid of chiplets integrated on an interposer, as illustrated in Fig. 1(a). Each chiplet is connected directly to the interposer via $\mu$-bumps surrounded by a capillary underfill material. The physical dimensions of the router ports are compatible with Universal Chiplet Interconnect Express (UCIe) specification (Sharma et al., 2022b). The entire package is covered by a copper lid, which contacts each chiplet through a thermal interface material (TIM).
@@
3D System Specifics: In the 3D system, three stacked chiplets are placed in a 4x4 grid with equal spacing, consistent with (Black et al., 2006). Vertically stacked chiplets are connected to one another with $\mu$-bumps surrounded by a capillary underfill material. The bottom chiplet of each stack is connected to the interposer with the same method. This integration is illustrated in Fig.  1(b). The lid contacts only the top chiplet layer through a thermal interface material.
5.2. Thermal RC Model Configuration
@@
The number of nodes in the thermal RC network determines the model complexity, runtime, and granularity at which temperature can be observed in the model. A higher node density is used in chiplets to optimize this trade-off, while fewer nodes are used in non-chiplet components, such as the interposer, lid, substrate, and so on. Each chiplet is divided into four equal quadrants. One node is placed within each quadrant to allow for granular temperature monitoring across each chiplet. An alternate node density is used for non-chiplet layers in each model, as described below. This easily configurable non-uniform node density enables higher thermal resolution in critical parts such as chiplets while decreasing the runtime with lower resolution in less critical structures such as the substrate and lid. The DSS models in our experimentation are created by discretizing the thermal RC models with $T_{s}=0.01s$ sampling period. The sampling time can be chosen according to the application requirements.
@@
3D Thermal RC Model Specifics: The node densities in the 3D system are adjusted similarly to the 2.5D models. The layers that contain chiplets use an 8$\times$8 grid, implying 4 nodes per chiplet. All other layers have a 4$\times$4 grid, leading to a lower node density. The entire 3D system consists of 48 chiplets in total. There are a total of 192 nodes counting all nodes within chiplets.
5.2.1. Input Workloads and Power Consumption
Identical workloads are considered for the 2.5D and 3D systems, with differences in chiplet power density detailed in the following subsections. The target chiplet systems are analyzed under one synthetic (WL1) and five real AI/ML application workloads (WL2-WL6). The synthetic workload starts with a stress test that applies the maximum power to all chiplets to increase temperature beyond 100C. Then, a pseudo-random bit sequence (PRBS) is applied to each chiplet to emulate a wide range of dynamic variations. Finally, all chiplets are turned off to let the temperature return to the ambient state, as depicted in Fig. 9. Besides testing transient and steady-state behaviors, this power profile helps us to tune the thermal RC model.
Table 8. Descriptions of the workloads used in this work. (C) and (I) denote the CIFAR100 and ImageNet datasets, respectively.
@@ 5.2.1. Input Workloads and Power Consumption
The remaining scenarios consider processing-in-memory (PIM)-based chiplets for accelerating ML workloads. The computational platform is resistive random access memory (ReRAM) based chiplets commonly used in literature (Sharma et al., 2022a; Shao et al., 2021; Song et al., 2017). We select this configuration due to its ability to efficiently implement matrix-vector multiplication, which is the predominant operation in any CNN workload. Each workload consists of a series of deep neural networks (DNNs) which run in series on the system. The workloads are listed in Table 8. The neural networks (NN) in these workloads consist of several networks such as ResNets, DenseNets, and VGG networks. For example, WL1 contains 16 ResNet34’s, followed by one VGG19, then 5 ResNet50’s, and so on. Each workload contains from 20 to 40 individual networks. Workloads are mapped to the system as computing resources become available, meaning a new NN is mapped to chiplets when it completes the execution of a previous NN. Consequently, these workloads consist of NNs ranging from small NNs like ResNet18, which can be mapped to a single chiplet, to larger NNs such as DenseNet169, which are spread across multiple chiplets.
After mapping the neural network workloads to the target chiplet-based system using a nearest-neighbor strategy inspired by the Simba (Shao et al., 2021), we estimate power consumption in two parts: computation and communication. Computation power is obtained using NeuroSim (Chen et al., 2018), while communication power is estimated using a custom version of BookSim that includes interconnect energy modeling (Jiang et al., 2013). We periodically sample the energy consumption from both tools at 10 ms intervals and convert these into dynamic power profiles for each chiplet. While these tools are not natively designed for chiplet systems, we follow a methodology adapted from SIAM that enables accurate per-chiplet power estimation for thermal simulations (Krishnan et al., 2021). Importantly, while we use this methodology for our experiments, MFIT only requires power traces as input, which can also be obtained from other simulators or real on-chip counters.
@@
Differences in 2.5D and 3D system chiplet power: Different hardware parameters, such as voltage and frequency, are used for the 2.5D and 3D systems. This results in lower per-chiplet power consumption in the 3D system of 1.2W as compared to 3W for the 2.5D system, detailed under the Power section of Table  7. Using these parameters, the total system power per lid area of the 3D system is between that of the 36 and 64 chiplet 2.5D system. This level means that the temperature of individual chiplets should be roughly equivalent between these systems, which is confirmed in Fig. 9.
5.2.2. Configuration of other thermal simulators: